for the function f(x) = x^4 - 2x^3 - 4x^2 + 4x - 1


-14 points BERRAPCALC7 2.5.001. For the function, find the following. f(x) = x4 – 2x3 – 4x2 + 4x - 1 (a) F(x) (b) f(x) (c)find f'(x), f''(x), f'''(x), f(4)(x)
Answer #3

f(x) = x -2 x² - 4x² + 4x - 1 f(x) = 4x² - 6x² - 8 x + 4 f(X) = 12x2-123-8 f(x) = 240-12 f(x) = 24
Answer #1
F'(x)=4x^3-6x^2-8x+4 F''(x)=12x^2-12x-8 F"'(x)=24x-12 F""(x)=24
source: 80
answered by: Margaret mercy
Answer #2
a) F'(x)=4x^3-6x^2-8x+4 b)F''(x)=12x^2-12x-8 C)F"'(x)=24x-12 d) F""(x)=24
source: 80
answered by: Margaret mercy